40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Mar 18 10:46:51 2010, Koji, Update, ABSL, PLL reconstructed 
    Reply  Thu Mar 18 12:27:56 2010, Alberto, Update, ABSL, PLL reconstructed 
       Reply  Thu Mar 18 13:40:35 2010, Koji, Update, ABSL, PLL reconstructed 
          Reply  Thu Mar 18 21:42:26 2010, Koji, Update, ABSL, PLL reconstructed 
             Reply  Mon Mar 22 22:11:26 2010, Mott, Update, ABSL, PLL reconstructed IMG_0150.JPG
                Reply  Mon Mar 22 23:37:32 2010, Mott, Update, ABSL, PLL reconstructed 2010-03-22_23.14.00.jpg2010-03-22_23.24.50.jpg
                   Reply  Tue Mar 23 18:44:46 2010, Mott, Update, ABSL, PLL reconstructed NPRO_PLL_freqresp.png
                      Reply  Wed Mar 24 03:58:18 2010, kiwamu, matt, koji, Update, Green Locking, leave PLL locked DSC_1402.JPG
                         Reply  Wed Mar 24 04:22:51 2010, kiwamu, matt, Update, Green Locking, two NPRO PLL  
                            Reply  Wed Mar 24 14:52:02 2010, Alberto, Update, Green Locking, two NPRO PLL  
                         Reply  Wed Mar 24 15:59:59 2010, kiwamu, matt, Update, Green Locking, leave PLL locked 
                            Reply  Thu Mar 25 17:29:48 2010, kiwamu, mott, Update, Green Locking, PLL two NPROs 
Message ID: 2696     Entry time: Mon Mar 22 22:11:26 2010     In reply to: 2684     Reply to this: 2697
Author: Mott 
Type: Update 
Category: ABSL 
Subject: PLL reconstructed 

 

 It looks like the PLL drifted alot over the weekend, and we couldn't get it back to 9 dBm.  We switched back to the new focus wideband PD to make it easier to find the beat signal.  I replaced all the electronics with the newly fixed UPDH box (#17) and we aligned it to the biggest beat frequency we could get, which ended up being -27 dBm with a -6.3V DC signal from the PD.  

Locking was still elusive, so we calculated the loop gain and found the UGF is about 45 kHz, which is too high.  We added a 20 dB attenuator to the RF input to suppress the gain and we think we may have locked at 0 gain.  I am going to add another attenuator (~6 dB) so that we can tune the gain using the gain knob on the UPDH box.  

Finally, attached is a picture of the cable that served as the smb - BNC adaptor for the DC output of the PD.  The signal was dependent on the angle of the cable into the scope or multimeter.  It has been destroyed so that it can never harm another innocent experiment again!

Attachment 1: IMG_0150.JPG  1.351 MB  | Hide | Hide all
IMG_0150.JPG
ELOG V3.1.3-